A chemical reaction involves breaking, making, or changing chemical bonds. These alterations in bonding lead to a change in eYou will determine a value for the heat capacity of your calorimeter by stud for which the heat of reaction is known: the forRun 1 Run 2 B Mass of weighing paper (g) Mass of weighing paper and Mg (p)00, 038隻0210-38629 Volume of HCl solution (mL) So.o0505050505050 0257025702570 r 55 5. 5. 6 6 6 6 7 7 7 7 8B Total mass of the final solution by adding the mass of initial HCI solution (density of 2 M HCl is 1.03 g/mL), the mass of

Ethalpy Change of a Chemical Reaction, Sample calculations for Run 2 B, all questions, please help me, thank you very much

A chemical reaction involves breaking, making, or changing chemical bonds. These alterations in bonding lead to a change in enthalpy, AHo, between the products and reactants. This enthalpy change takes into account the change in intermal energy in the system and the work done by the system on its surroundings during the reaction. If the reaction occurs at a constant pressure and in an adiabatic apparatus (no heat can leave or enter), the change in enthalpy will be equal to the heat absorbed or evolved, reflected by the temperature of the system decreasing or increasing, respectively. The heat of a reaction is determined in the laboratory by measuring this temperature change in a calorimeter. If you know the amount of heat required to change the temperature of the calorimeter, reactants, and products, you can calculate the heat of reaction or the enthalpy change In this experiment you will measure the heat evolved in the reaction of magnesium with hydrochloric acid Mg(s)+2 H (a)- Mg2 (aą)+H2 (g) In order to calculate the heat of this reaction very precisely, you would need to know the heat capacities (heat required to change the temperature of one gram one degree Celsius) of everything that changes temperature during the reaction, and the magnitudes of corrections for the amount of heat lost to the surroundings and for the heat produced by the mechanical stirring of the contents of the calorimeter. To simplify the calculations in this experiment, we will assume that the reactants will not change temperature during the reaction, as they will be consumed quickly, and that the heat absorbed by the hydrogen gas formed will be too small to add appreciably to the other terms. The corrections for heat loss and the other factors related to the apparatus will be included in one term, which ill call the heat capacity of the calorimeter
You will determine a value for the heat capacity of your calorimeter by stud for which the heat of reaction is known: the formation of water from its ions solution in dil ut H (a)+OH (a4) HO(a) 7,320 Joules
Run 1 Run 2 B Mass of weighing paper (g) Mass of weighing paper and Mg (p)00, 038隻0210-38629 Volume of HCl solution (mL) So.om Co.oml Molarity of HCl solution (M) Volume of water added to HCI solution (mL) Initial temperature of HCl solution (ec) Temperatures of solution (°C) at 0.25- min (15-sec) intervals 리,81 0.50 1.00 1.50 2.00 2.50 3.00 3.50 4.00 4.25 4.50 4.75 5.00 2 22,0°C 220°C | 240°C
0505050505050 0257025702570 r 55 5. 5. 6 6 6 6 7 7 7 7 8
B Total mass of the final solution by adding the mass of initial HCI solution (density of 2 M HCl is 1.03 g/mL), the mass of water (density of water is 1.00 g/mL), and the mass of magnesium Heat required to raise the MgCI2 solution to the maximum temperature assuming that the heat capacity of your final MgCl solution is 3.97 J/g "C (from PRE-LABORATORY QUESTION 5) Heat required to raise the calorimeter to the maximum temperature using the value for the heat capacity of the calorimeter as determined in part A of this experiment (from PRE-LABORATORY QUESTION 6) Total heat evolved by the reaction by adding the heats required to raise both the solution and the calorimeter to the maximum temperature Heat evolved per mole of Mg Change in enthalpy for the following specific reaction. Be sure to indicate the correct sign (- or+) for the change in enthalpy Mg(s) +2H+ (aq) → Mg" (aq) + H2 (g)
Answer #1

The Total mass of the solution is

= Mass of water + Mass of 2.0 M of HCl + Mass of Mg added

Mass of water = volume of water x desity of water = 50 x 1 = 50g

Mass of HCl = Volume of HCl x Desity of HCl = 2 x 1.03 = 2.06

Mass of Mg added = 0.8863

Total mass = 50 + 2.06 + 0.8863 = 54.9463 g

******************************

The heat required to raise to the temperature for final MgCl solution is

= \DeltaT x Mass of MgCl solution x specific heat capacity of MgCl solution

\DeltaT = maximum Temperature(T2) - Temperature before reaction (T1)

T1 = 21 0C and T2 = 24.0 0C

\DeltaT = 3 oC , Mass of solution = 54.9463 g and Specific heat capacity of MgCl = 3.97 J/g oC

So, heat required to raise to the temperature for final MgCl solution is

= 21 x 54.9463 x 3.97 = 4580 J

******************

The heat required to raise calorimeter to maximum temperature

It has to be caliculated. It will differ from one to another. Let us assume it will be 1500 J

**********************

The total heat evolved per 0.8863 g of Mg reaction with HCl is

= Heat required to raise MgCl solution + Heat required to raise calorimeter to a maximum temperature

= 4580 J + 1500 J = 6080 J

****************

The total heat evolved per mole of Mg is

= Total heat evolved / no moles of Mg used

Hear no of moles of Mg used is

= wt of Mg / atomic wt of Mg = 0.8863 / 24.30 = 0.0364

The total heat evolved per mole of Mg is

= 6080 / 0.0364 = 167033 J / Mol = 16.7033 KJ /Mol

**********************************

\DeltaH0 = -167033 J /Mol (The negative sign indicate that heat is releasing from reaction, So it is a exothermic reaction)

for

Mg + 2H+ ----------------> Mg2+ + H2